Which is the standard form of the eqquation of the line that is perpendicular to 8x+6y=9 and contains (2,1)
Answer by lwsshak3(11628)

Which is the standard form of the eqquation of the line that is perpendicular to Which is the standard form of the eqquation of the line that is perpendicular to 8x+6y=9 and contains (2,1)
**
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept.
8x+6y=9
rewrite in standard form:
6y=-8x+9
y=-(8/6)x+9/6
y=-(4/3)x+3/2
slope=-4/3
...
For perpendicular line:
slope=3/4 (negative reciprocal of line above)
Equation: y=(3/4)x+b
solving for b using coordinates of given point (2,1)
1=(3/4)*2+b
b=1-6/4=-1/2
Equation of perpendicular line: y=(3/4)x-1/2